OpenStudy (anonymous):

after finding the volume of a cone, how is it written if the exact volume in terms of pi is being asked for?

OpenStudy (amistre64):

let pi be pi instead of 3.14 of some other thing

OpenStudy (amistre64):

\[V=\frac13\pi r^2h\]just let pi be pi and solve the scalar:\[\frac{r^2h}{3}\pi\]

OpenStudy (amistre64):

\[\frac{r^2h}{3}\pi\approx\frac{r^2h}{3}(3.1415...)\]